In a town every person takes at least one cup of coffee or tea. if 3/5th of the population takes coffee, 14/25th of the population takes tea, then the percentage of population which takes both coffee and tea is??
a)16 b)15 c)18) d)20

Answer by flame8855(424)

assume a population of 25
so 3/5 takes coffee so 3/5*25 = 15 takes coffe
so 25- 15 = 10 people only tea
14/25 * 25= 14 people tea so
25-14= 11 people coffee only
11+10 = 21
25-21= 4 people coffee and tea both
4/25*100 = 16 %
so the answer is a